# ICon: In-Context Contribution for Automatic Data Selection

### v1

@@ 3.3 ICon-guided Selection Paradigm Training
ICon (5%) 16.21 28.96 56.91 27.27 53.70 40.51 47.18 44.23 58.44 24.42 37.75 69.46 42.09(+4.14) 3.06
ICon (10%) 16.08 27.99 57.86 26.76 53.57 42.63 45.48 42.81 56.27 26.57 42.01 66.93 42.08(+4.13) 5.62
ICon (15%) 21.16 29.89 60.58 30.81 52.30 43.16 44.80 44.15 55.56 28.42 41.64 67.96 43.37(+5.42) 8.37
@@
ICon (5%) 20.85 36.25 68.34 27.88 64.73 43.26 53.50 47.80 58.10 32.10 39.72 69.06 46.80(+0.15) 2.27
ICon (10%) 25.28 37.05 69.26 34.85 63.67 43.35 50.00 47.00 57.85 35.79 40.25 68.27 47.72(+1.07) 5.28
ICon (15%) 23.07 30.11 70.46 34.34 64.09 46.90 51.71 47.68 58.16 38.40 41.96 67.80 47.89(+1.24) 6.85
@@
ICon (5%) 7.78 31.47 51.93 23.23 36.95 45.72 40.53 36.07 57.57 23.96 40.62 66.77 38.55(+0.01) 3.26
ICon (10%) 11.86 29.76 50.62 25.76 37.30 45.19 41.72 35.38 57.39 24.88 39.82 66.38 38.84(+0.30) 6.16
ICon (15%) 8.06 26.83 54.67 27.78 36.39 43.89 42.66 34.74 57.01 23.50 39.80 66.06 38.45(-0.09) 9.18
Table 1: Performance of LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B in terms of FLOPs and evaluation benchmarks for instruction following, knowledge, and reasoning. The datasets TQA, HS, LQA, and WG correspond to TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. Models trained on ICon-selected data outperform those trained on full datasets on evaluation benchmarks, achieving better performance with fewer samples.
@@ 4.1 Datasets
In this paper, we use the classic Alpaca dataset Taori et al. (2023), one of the original instruction datasets created by Stanford University. The Alpaca dataset consists of 52,002 instruction-following samples, generated with TextDavinci-003 through the self-instruction approach Wang et al. (2022). The LLM trained on this dataset demonstrate a good instruction-following capability. However, its heavy dependence on TextDavinci-003 has raised concerns among researchers regarding the quality of instruction data. The WizardLM dataset Xu et al. (2023) utilizes the Evol-Instruct algorithm to enhance the quality and complexity of instruction data. We conduct experiments using WizardLM-70K to verify the generalization of our method across different datasets.
@@
We consider the impact of a training candidate sample on the model’s performance over the assessment dataset as an indicator of its overall influence on model capability, as constructed in Section 3.1. To reduce bias and ensure diversity, we sample data from three sources: OpenOrca-GPT3.5 Lian et al. (2023) containing ChatGPT-generated data, OpenOrca-GPT4 containing GPT-4-generated data, and Dolly-15K Mike et al. (2023) containing human-generated data. From these three datasets, we randomly select 1,020 instructions to construct the assessment dataset.
@@ 4.2 Implementation Details
For experiments on the LLaMA3.1-8B Grattafiori et al. (2024), Qwen2.5-3B Yang et al. (2024), and LLaMA2-7B Lyu et al. (2024) models, we all follow the original Alpaca [^2] https://github.com/tatsu-lab/stanford_alpaca training configuration using the Alpaca codebase. Models are trained for three epochs with the Adam optimizer Kingma and Ba (2014), a learning rate of $2\times 10^{-5}$, and a batch size of 128. The maximum input length is set to 512 for Alpaca dataset and 2048 for WizardLM dataset.
@@ 4.3.1 Benchmark Evaluation
To evaluate the effectiveness of ICon, we report performance on several widely used benchmarks. These benchmarks fall into three categories, assessing different capabilities: instruction following, knowledge, and reasoning. Instruction following evaluation includes two key benchmarks: IFEval Zhou et al. (2023c) and AlpacaEval Li et al. (2023b). AlpacaEval provides an LLM-based automatic evaluation on the AlpacaFarm set Dubois et al. (2023), where we use GPT-4 Achiam et al. (2023) as both the response generator and the evaluator comparing model outputs with reference responses. Knowledge evaluation includes four key benchmarks: GLUE Wang et al. (2018), GPQA Rein et al. (2024), MMLU Hendrycks et al. (2020), and TruthfulQA Lin et al. (2021). Reasoning evaluation includes six key benchmarks: ARC Clark et al. (2018), BBH Suzgun et al. (2022), HellaSwag Zellers et al. (2019), LogiQA Liu et al. (2020), MuSR Sprague et al. (2023), and Winogrande Sakaguchi et al. (2021).
@@ 4.3.2 Pairwise Comparison
The performance on several pairwise comparison test sets for instruction following is also provided, including WizardLM Xu et al. (2023), Self-Instruct Wang et al. (2022), Vicuna Chiang et al. (2023), Koala Vu et al. (2023), and LIMA Zhou et al. (2023b). These sets contain 218, 252, 80 and 300 human-curated instruction data, respectively, across domains such as math, coding, writing, and general knowledge. We use GPT-4 Hurst et al. (2024) as the judge. For each instruction in the test sets, candidate models generate responses, which are then scored by GPT-4 in a pairwise fashion. The judge rates each response from 1 to 10 based on factors such as relevance and accuracy. To mitigate positional bias Wang et al. (2023a), we submit each response pair twice with reversed order. A model is considered to win only if it does not lose in both orders, following these rules: Win: wins both, or wins one and ties the other; Tie: ties both, or wins one and loses one; Lose: loses both, or loses one and ties the other.
@@ 5.1 Main Results
In this section, we present the evaluation benchmark results shown in Table 1. The models are trained using different proportions of Alpaca data selected by the ICon method, including 1%, 5%, 10%, and 15%, corresponding to 520, 2,600, 5,200, and 7,800 samples, respectively. "All" denotes models trained on the full, unfiltered dataset. Experiments were conducted on multiple models, including LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B. Across all benchmarks, models trained on a small fraction of ICon-selected data often match or surpass the performance of those trained on the full dataset, and consistently outperform them in overall metrics. For example, the LLaMA3.1-8B model trained on just 15% of ICon-selected data exceeds the full-data baseline by 5.42 percentage points. Similarly, the Qwen2.5-3B model with 15% ICon-selected data and the LLaMA2-7B model with only 1% achieve overall gains of 1.24% and 0.65%, respectively. Notably, these ICon-trained models require only about 1/5 to 1/100 of the FLOPs compared to full-data training, demonstrating significantly lower resource consumption. These results underscore the effectiveness of ICon-guided selection in enhancing instruction tuning with substantially less data.
The results of the pairwise evaluation, including detailed win-tie-lose statistics on the Vicuna, Koala, WizardLM, SInstruct, and LIMA test sets, are presented in Figure 2. We focus on the best-performing models from the evaluation benchmarks: LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B, each trained on 7,800, 7,800, and 520 ICon-selected samples, respectively. All three models consistently outperform their counterparts trained on the full Alpaca dataset across all five test sets. These results further validate the effectiveness of ICon-guided selection in improving instruction-following ability, even with significantly less training data.
5.2 Comparison with Other Methods
Comparison Pairwise Winning Score $\uparrow$ Overall
Test Set Vicuna WizardLM LIMA SInstruct Koala
ICon vs. Random 1.2125 1.4541 1.3233 1.5159 1.3389 1.3922
ICon vs. Low PPL 1.3375 1.3945 1.4533 1.3690 1.3333 1.3903
ICon vs. Top PPL 1.9126 1.9125 1.8567 1.8452 1.7778 1.8556
ICon vs. Alpagasus Chen et al. (2023) 1.1250 1.1789 1.1033 1.3611 1.3389 1.2252
ICon vs. Deita Liu et al. (2023c) 1.3375 1.2661 1.2733 1.3730 1.3889 1.3214
ICon vs. Superfilter Li et al. (2024b) 1.0375 1.0780 1.0100 1.1627 1.0944 1.0786
Table 2: Comparison with other methods on Vicuna, WizardLM, LIMA, SInstruct, and Koala test set. The pairwise Winning Scores are calculated between models using our method and other methods. All the comparisons are performed by GPT-4, and the values that are greater than 1.0 represent our models are better and vice versa. The models are trained on LLaMA3.1-8B with 7,800 selected samples. ICon consistently outperforms widely used data selection methods on pairwise comparison test sets.
In this subsection, we compare our method with several widely used data selection baselines on the Alpaca dataset, as shown in Table 2 and Table 4. "Random" selects samples at random. "Low PPL" selects samples with the lowest perplexity, while "Top PPL" selects those with the highest perplexity. "Alpagasus" uses GPT-3.5-Turbo to score sample responses based on dimensions such as helpfulness and accuracy. "Deita" uses a set of models fine-tuned from LLaMA and Mistral to evaluate sample quality, complexity, and diversity. "Superfilter" selects samples using IFD scores that reflect instruction difficulty, computed using GPT-2. Superfilter assumes that IFD scores from smaller models are as reliable as those from larger ones.
[…]
6 Analysis
Optimal Data Scale for Instruction Tuning
We further investigate how the scale of selected data affects model performance. Using the LLaMA3.1-8B model as a case study, Figure 3 presents the results of training on ICon-selected subsets of the Alpaca dataset at varying proportions: 1%, 5%, 10%, 15%, 20%, 25%, 30%, 50%, 75%, and 100%. The average score across all evaluation benchmarks generally rises and then declines, peaking at 15%, demonstrating that a small, carefully selected subset can outperform models trained on much larger datasets. Based on this observation, we select the model trained on 15% (7,800 samples) of ICon-selected data as the representative model for subsequent comparisons. Detailed results for each selection scale are provided in Appendix B. These results show that adding more data beyond 15% does not lead to further improvements, highlighting the efficiency of ICon in identifying the most contributive samples.
Figure 3: Model performance (average score) on evaluation benchmarks with varying proportions of ICon-selected Alpaca data. The models are trained based on LLaMA3.1-8B. The dataset is Alpaca dataset. As data scale grows, performance generally improves and then declines, with the best result at 15%.
Generalization of ICon on different Datasets
@@ 6 Analysis
FULL 33.61 45.46 44.37 42.94
[…]
ICon (5%) 43.21 48.28 48.36 47.48
ICon (10%) 36.89 48.15 47.18 45.79
ICon (15%) 35.24 46.29 45.76 44.18
Table 3: Evaluation benchmark results of models trained on WizardLM dataset and its ICon-selected subsets. The ‘IF’, ‘KN’, and ‘RS’ correspond to average scores on Instruction Following, Knowledge, and Reasoning, respectively. Results on WizardLM dataset confirm the generalization of ICon across datasets.
IF KN RS Avg
@@
Table 4: Performance of different methods on evaluation benchmarks. The ‘IF’, ‘KN’, and ‘RS’ correspond to average scores on Instruction Following, Knowledge, and Reasoning, respectively. The models are trained on LLaMA3.1-8B with 7,800 selected samples. ICon outperforms widely used data selection methods on pairwise evaluation benchmarks.
@@
In addition to the main experiments conducted on the Alpaca dataset, we also evaluated the generalization of the ICon method to different datasets. Specifically, we performed experiments on the WizardLM dataset. The results, shown in Table 6 and detailed in Appendix C, cover selection scales of 1%, 5%, 10%, 15%, and ALL, corresponding to 520, 2,600, 5,200, 7,800, and 70K samples. As shown, models trained on ICon-selected subsets consistently outperform those trained on the full dataset across the Instruction Following, Knowledge, and Reasoning tasks, achieving higher average scores and overall performance. The model trained on just 5% of the data achieves the best results, with a 4.54 percentage point improvement in average evaluation benchmark score. These findings demonstrate the effectiveness of ICon across datasets. Notably, the selection paradigm, trained on the Alpaca subset, is reused here without modification, demonstrating its robust generalization.
Characteristics of High-Contribution ICon Samples
@@
Figure 4: Visualization using t-SNE on sample embeddings from the Alpaca dataset. Red points represent samples with the top 15% ICon high-contribution scores and gray points represent other samples from the dataset. High-contribution ICon samples exhibit diverse characteristics.
@@
Figure 5: Difficulty distribution of top 15% high-contribution samples vs. full Alpaca dataset. High-contribution samples (red bars), selected via ICon for instruction tuning, exhibit distinct difficulty patterns compared to the full dataset (gray bars). High-contribution ICon samples tend to fall within an appropriate difficulty range.
@@
We further analyzed the characteristics of the high-contribution samples selected by the ICon method. Figure 4 shows the distribution of high-contribution samples in the t-SNE space, where we observe both cross-cluster dispersion and local density. This pattern aligns with the notion in previous studies Liu et al. (2023c); Bukharin and Zhao (2023) that "instruction data should be diverse". However, the distribution is not uniformly scattered. Instead, some clusters exhibit higher density, suggesting that high-contribution samples tend to exhibit certain clustering behavior in the feature space. Figure 5 further compares the distribution of instruction difficulty between high-contribution samples and the full dataset. We use the IFD score Li et al. (2023a, 2024b) as the difficulty metric. The t-test Kendall (1937) result (t = 2.24, p = 0.0251) shows that high-contribution samples are significantly more difficult than the full dataset on average, indicating that the selection method tends to favor more challenging samples. However, as seen in the distribution, high-contribution samples are not exclusively concentrated in the highest difficulty range. Interestingly, while the method favors more difficult samples, some mid-level difficulty samples are retained, and extreme difficulty samples are avoided. This finding challenges the view that more difficult samples are always inherently more beneficial for model training Du et al. (2023); Li et al. (2023a, 2024b, 2024a); Zhang et al. (2025).

### v2

@@ 3.3 RICo-guided Selection Paradigm
RICo (5%) 16.21 28.96 56.91 27.27 53.70 40.51 47.18 44.23 58.44 24.42 37.75 69.46 42.09(+4.14) 3.06
RICo (10%) 16.08 27.99 57.86 26.76 53.57 42.63 45.48 42.81 56.27 26.57 42.01 66.93 42.08(+4.13) 5.62
RICo (15%) 21.16 29.89 60.58 30.81 52.30 43.16 44.80 44.15 55.56 28.42 41.64 67.96 43.37(+5.42) 8.37
@@
RICo (5%) 20.85 36.25 68.34 27.88 64.73 43.26 53.50 47.80 58.10 32.10 39.72 69.06 46.80(+0.15) 2.27
RICo (10%) 25.28 37.05 69.26 34.85 63.67 43.35 50.00 47.00 57.85 35.79 40.25 68.27 47.72(+1.07) 5.28
RICo (15%) 23.07 30.11 70.46 34.34 64.09 46.90 51.71 47.68 58.16 38.40 41.96 67.80 47.89(+1.24) 6.85
@@
RICo (5%) 7.78 31.47 51.93 23.23 36.95 45.72 40.53 36.07 57.57 23.96 40.62 66.77 38.55(+0.01) 3.26
RICo (10%) 11.86 29.76 50.62 25.76 37.30 45.19 41.72 35.38 57.39 24.88 39.82 66.38 38.84(+0.30) 6.16
RICo (15%) 8.06 26.83 54.67 27.78 36.39 43.89 42.66 34.74 57.01 23.50 39.80 66.06 38.45(-0.09) 9.18
Table 1: Performance of LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B in terms of FLOPs and evaluation benchmarks for instruction following, knowledge, and reasoning. The datasets TQA, HS, LQA, and WG correspond to TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. Models trained on RICo-selected data outperform those trained on full datasets on evaluation benchmarks, achieving better performance with fewer samples.
@@ 4.1 Datasets
In this paper, we use the classic Alpaca dataset Taori et al. (2023), one of the original instruction datasets created by Stanford University. The Alpaca dataset consists of 52,002 instruction-following samples, generated with TextDavinci-003 through the self-instruction approach Wang et al. (2022b). The WizardLM dataset Xu et al. (2023) utilizes the Evol-Instruct algorithm to enhance the quality and complexity of instruction data. We conduct experiments using WizardLM-70K to verify the generalization of our method across different datasets.
@@
We consider the impact of a training candidate sample on the model’s performance over the assessment dataset as an indicator of its overall influence on model capability, as constructed in Section 3.1. To reduce bias and ensure diversity, we sample data from three sources: OpenOrca-GPT3.5 Lian et al. (2023) containing ChatGPT-generated data, OpenOrca-GPT4 containing GPT-4-generated data, and Dolly-15K Mike et al. (2023) containing human-generated data. From these three datasets, we randomly select 1,020 instructions to construct the assessment set.
@@ 4.2 Implementation Details
For experiments on the LLaMA3.1-8B Grattafiori et al. (2024), Qwen2.5-3B Yang et al. (2024), and LLaMA2-7B Lyu et al. (2024), we all follow the original Alpaca [^2] https://github.com/tatsu-lab/stanford_alpaca training configuration using its codebase. Models are trained for three epochs with Adam optimizer Kingma and Ba (2014), a learning rate of $2\times 10^{-5}$, and a batch size of 128. The maximum input length is set to 512 for Alpaca dataset and 2048 for WizardLM dataset.
@@ 4.3.1 Benchmark Evaluation
To evaluate the effectiveness of RICo, we report performance on several widely used benchmarks. These benchmarks fall into three categories, assessing different capabilities: instruction following, knowledge, and reasoning. Instruction following evaluation includes: IFEval Zhou et al. (2023c) and AlpacaEval Li et al. (2023b). AlpacaEval provides automatic evaluation on the AlpacaFarm set Dubois et al. (2023), where we use GPT-4 Achiam et al. (2023) as both the response generator and evaluator. Knowledge evaluation includes: GLUE Wang et al. (2018), GPQA Rein et al. (2024), MMLU Hendrycks et al. (2020), and TruthfulQA Lin et al. (2021). Reasoning evaluation includes: ARC Clark et al. (2018), BBH Suzgun et al. (2022), HellaSwag Zellers et al. (2019), LogiQA Liu et al. (2020), MuSR Sprague et al. (2023), and Winogrande Sakaguchi et al. (2021).
@@ 4.3.2 Pairwise Comparison
The performance on pairwise comparison test sets for instruction following is also provided, including WizardLM Xu et al. (2023), Self-Instruct (SInstruct)  Wang et al. (2022b), Vicuna Chiang et al. (2023), Koala Vu et al. (2023), and LIMA Zhou et al. (2023b). These sets contain 218, 252, 80 and 300 human-curated instruction data, respectively, across domains such as math, coding, writing, and general knowledge. We use GPT-4 Hurst et al. (2024) as the judge. For each instruction, candidate models generate responses, which are scored by GPT-4 in a pairwise fashion on a scale of 1 to 10 based on relevance, accuracy, and other criteria. To mitigate positional bias Wang et al. (2023a), we submit each response pair twice with reversed order. A model is considered to win only if it does not lose in both orders, following these rules: Win: wins both, or wins one and ties the other; Tie: ties both, or wins one and loses one; Lose: loses both, or loses one and ties the other.
@@ 5.1 Main Results
In this section, we present the evaluation benchmark results shown in Table 1. The models are trained using different proportions of Alpaca data selected by the RICo method, including 1%, 5%, 10%, and 15%, corresponding to 520, 2,600, 5,200, and 7,800 samples, respectively. FULL denotes models trained on the full, unfiltered dataset. Experiments are conducted on multiple models, including LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B. Models trained on a small fraction of RICo-selected data often match or exceed the performance of full-data models across all benchmarks, with consistent gains in average benchmark scores. Specifically, the LLaMA3.1-8B model trained on just 15% of RICo-selected data exceeds the full-data baseline by 5.42 percentage points. Similarly, the Qwen2.5-3B model with 15% RICo-selected data and the LLaMA2-7B model with only 1% achieve overall gains of 1.24% and 0.65%, respectively. Notably, these RICo-trained models require only about 1/5 to 1/100 of the FLOPs compared to full-data training, demonstrating significantly lower resource consumption. These results underscore the effectiveness of RICo in enhancing instruction tuning with substantially less data.
@@
Random 12.13 45.18 46.89 40.52
@@
Table 2: Performance of different methods on evaluation benchmarks. The ‘IF’, ‘KN’, and ‘RS’ correspond to average scores on Instruction Following, Knowledge, and Reasoning, respectively. The models are trained on LLaMA3.1-8B with 15% selected samples. RICo outperforms widely used data selection methods on evaluation benchmarks.
The results of the pairwise evaluation, including detailed win-tie-lose statistics on the Vicuna, Koala, WizardLM, SInstruct, and LIMA test sets, are presented in Figure 2. We focus on the best-performing models from the evaluation benchmarks: LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B, each trained on 15%, 15%, and 1% RICo-selected samples, respectively. All three models consistently outperform their counterparts trained on the full Alpaca dataset across all five test sets. These results further validate the effectiveness of RICo in improving instruction-following ability, even with significantly less training data.
5.2 Comparison with Other Methods
Comparison Pairwise Winning Score $\uparrow$ Overall
Test Set Vicuna WizardLM LIMA SInstruct Koala
RICo vs. Random 1.2125 1.4541 1.3233 1.5159 1.3389 1.3922
RICo vs. Low PPL 1.3375 1.3945 1.4533 1.3690 1.3333 1.3903
RICo vs. Top PPL 1.9125 1.8073 1.8567 1.8452 1.7778 1.8340
RICo vs. Alpagasus Chen et al. (2023) 1.1250 1.1789 1.1033 1.3611 1.3389 1.2252
RICo vs. Deita Liu et al. (2023c) 1.3375 1.2661 1.2733 1.3730 1.3889 1.3214
RICo vs. Superfilter Li et al. (2024b) 1.0375 1.0780 1.0100 1.1627 1.0944 1.0786
RICo vs. Nuggets Li et al. (2023c) 1.1625 1.1422 1.0967 1.1508 1.1111 1.1272
Table 3: Comparison with other methods on Vicuna, WizardLM, LIMA, SInstruct, and Koala test set. The pairwise Winning Scores are calculated between models using our method and other methods. All the comparisons are performed by GPT-4, and the values that are greater than 1.0 represent our models are better and vice versa. The models are trained on LLaMA3.1-8B with 15% selected samples. RICo consistently outperforms widely used data selection methods on pairwise comparison test sets.
In this section, we compare our method with several widely used data selection baselines on the Alpaca dataset, as shown in Table 3 and Table 2. Random selects samples uniformly at random. Low-PPL and Top-PPL select samples with the lowest and highest perplexity scores, respectively. Alpagasus Chen et al. (2023) uses GPT-3.5-Turbo to score responses based on helpfulness, accuracy, and other dimensions. Deita Liu et al. (2023c) leverages models fine-tuned from LLaMA and Mistral to assess sample quality, complexity, and diversity. Superfilter Li et al. (2024b) ranks samples using IFD scores derived from GPT-2 to estimate instruction difficulty, assuming that smaller models provide sufficiently reliable signals. Nuggets Li et al. (2023c), a representative ICP-based method, uses binary comparisons in the ICL setting to assign golden scores. We adopt its KMeans100 selection method, consistent with its official implementation.
[…]
6 Analysis
Optimal Data Scale for Instruction Tuning
We further investigate how the scale of selected data affects model performance. Using the LLaMA3.1-8B model as a case study, Figure 3 presents the results of training on RICo-selected subsets of the Alpaca dataset at varying proportions: 1%, 5%, 10%, 15%, 20%, 25%, 30%, 50%, 75%, and 100%. The average score across all evaluation benchmarks generally rises and then declines, peaking at 15%, indicating that there exists an optimal data scale for instruction tuning. The performance drop beyond this point suggests that low-contribution samples begin to dilute the training signal, underscoring the effectiveness of RICo in prioritizing high-contribution data. Based on this observation, we select the model trained on 15% of RICo-selected data as the representative model for comparisons. Detailed results for each selection scale are provided in Appendix B.
Figure 3: Model performance (average score) on evaluation benchmarks with varying proportions of RICo-selected Alpaca data. The models are trained based on LLaMA3.1-8B. The dataset is Alpaca dataset. As data scale grows, performance generally improves and then declines, with the best result at 15%.
Generalization of RICo on different Datasets
IF KN RS Avg
FULL 33.61 45.46 44.37 42.94
[…]
RICo (5%) 43.21 48.28 48.36 47.48
RICo (10%) 36.89 48.15 47.18 45.79
RICo (15%) 35.24 46.29 45.76 44.18
Table 4: Evaluation benchmark results of models trained on WizardLM dataset and its RICo-selected subsets. The ‘IF’, ‘KN’, and ‘RS’ correspond to average scores on Instruction Following, Knowledge, and Reasoning, respectively. Results on WizardLM dataset confirm the generalization of RICo across datasets.
@@ 6 Analysis
In addition to the main experiments on the Alpaca dataset, we evaluate the generalization of RICo to a different dataset. Specifically, we apply the selection paradigm trained on Alpaca samples to directly filter the WizardLM dataset. The results, shown in Table 6 and detailed in Appendix C, cover selection scales of 1%, 5%, 10%, 15%, and 100%, corresponding to 700, 3,500, 7,000, 10,500, and 70K samples, respectively. As shown, models trained on RICo-selected subsets consistently outperform those trained on the full dataset across Instruction Following, Knowledge, and Reasoning tasks, achieving higher average and overall scores. Notably, the model trained on just 5% of the data achieves the best results, with a 4.54 percentage point improvement in average evaluation benchmark score. These findings highlight the effectiveness of RICo across datasets. Importantly, the selection paradigm used here is directly transferred from Alpaca without modification, demonstrating its robust cross-dataset generalization.
Characteristics of High-Contribution RICo Samples
@@
Figure 4: Visualization using t-SNE on sample embeddings from the Alpaca dataset. Red points represent samples with the top 15% RICo high-contribution scores and gray points represent other samples from the dataset. High-contribution RICo samples exhibit diverse characteristics.
@@
Figure 5: Difficulty distribution of top 15% high-contribution samples vs. full Alpaca dataset. High-contribution samples (red bars), selected via RICo for instruction tuning, exhibit distinct difficulty patterns compared to the full dataset (gray bars). High-contribution RICo samples tend to fall within an appropriate difficulty range.
@@
We further analyzed the characteristics of the high-contribution samples selected by the RICo method. Figure 4 shows the distribution of high-contribution samples in the t-SNE space, where we observe both cross-cluster dispersion and local density. This pattern aligns with the notion in previous studies Liu et al. (2023c); Bukharin and Zhao (2023) that "instruction data should be diverse". However, the distribution is not uniformly scattered. Instead, some clusters exhibit higher density, suggesting that high-contribution samples tend to exhibit certain clustering behavior in the feature space. Figure 5 further compares the distribution of instruction difficulty between high-contribution samples and the full dataset. We use the IFD score Li et al. (2023a, 2024b) as the difficulty metric. The t-test Kendall (1937) result (t = 2.24, p = 0.0251) shows that high-contribution samples are significantly more difficult than the full dataset on average, indicating that the selection method tends to favor more challenging samples. However, as seen in the distribution, high-contribution samples are not exclusively concentrated in the highest difficulty range. Interestingly, while the method favors more difficult samples, some mid-level difficulty samples are retained, and extreme difficulty samples are avoided. This finding challenges the view that more difficult samples are always inherently more beneficial for model training Du et al. (2023); Li et al. (2023a, 2024b, 2024a); Zhang et al. (2025).